# Structure-Dynamics Correlation and Its Link to Fragility and Dynamic Heterogeneity

### v2

@@ Abstract
Understanding the connection between structure, dynamics, and fragility (the rate at which relaxation time grow with decreasing temperature) is central to unravelling the glass transition. Fragility is often linked to dynamic heterogeneity, and thus it is commonly assumed that if structure influences dynamics, more fragile systems should exhibit stronger structure–dynamics correlations. In this study, we test the generality of this assumption using three model systems: Lennard-Jones (LJ) and Weeks–Chandler–Andersen, where fragility is tuned via density, and a modified LJ $(q,p)$ system, where potential softness is changed to vary fragility. We employ a structural order parameter derived from the mean-field caging potential and analyse energy barriers at both macroscopic and microscopic levels. While the macroscopic slope of the energy barrier, suitably defined, correlates with fragility, no consistent correlation is found for the microscopic energy barriers. Instead, the latter shows a strong correlation with an independently computed structure–dynamics measure obtained from isoconfigurational ensemble. Surprisingly, the two systems with the highest structure–dynamics correlation, LJ at $\rho=1.1$ and the $(8,5)$ model, are respectively the least and most fragile within their classes. These systems exhibit broad mobility distributions, bimodal displacement profiles, and high non-Gaussian parameters, all indicative of dynamic heterogeneity. However, their dynamic susceptibilities remain low, suggesting a decoupling between spatial correlation and temporal heterogeneity. Both systems lie in the enthalpy-dominated regime and are near the spinodal, suggesting mechanical instability as a source of heterogeneity. These findings challenge the conventional linkage among fragility, heterogeneity, and structure–dynamics correlation.
@@ I Introduction
The behaviour of supercooled liquids near the glass transition is characterised by a dramatic slowing down in dynamics without any obvious structural transition Debenedetti2001_Nature ; Ediger2000_Physchem ; Dyre2006_modern ; Lubchenko2007_Physchem ; Tong2018_prx . The rapidity with which the relaxation time increases with decreasing temperature is quantified in terms of the fragility parameter Angell1995_science . Fragile liquids show a super-Arrhenius increase in relaxation time, while strong liquids follow an Arrhenius behaviour Martinez2001_nature .
@@
One of the hallmarks of supercooled liquids is dynamic heterogeneity, the presence of regions with significantly different mobilities. Traditionally, this heterogeneity has been linked to fragility: more fragile systems tend to exhibit stronger dynamic heterogeneity Martinez2001_nature ; Ediger2000_Physchem ; Sillescu1999_noncryst ; Donati1998_pre ; Kob1997_prl ; Biroli2006_prl ; Sastry2001_nature . This link is often assessed through quantities like the four-point dynamic susceptibility, $\chi_{4}$, which captures spatial correlations in particle dynamics Lacevic2003_jcp ; Toninelli2005_pre .
@@
Although, as discussed in the literature, the glass transition and the rapid growth of the dynamics happens without substantial structural change Debenedetti2001_Nature ; Lubchenko2007_Physchem , studies have shown that structure does indeed contribute to the dynamics, and small changes in structure can lead to large changes in other quantities like entropy which eventually strongly effects the dynamics Banerjee2014_prl ; Banerjee2016_jcp ; Nandi2015_jcp . There have also been studies trying to understand the structural origin of dynamical heterogeneity Cooper2008_nature ; Schoenholz2015_nature ; Nandi2021_prl ; Richard2020_prm . Among these studies, the machine learning study expressing the structural order by a softness parameter Schoenholz2015_nature and the microscopic study involving some of us expressing the structural order in terms of the depth of a mean field caging potential Nandi2021_prl ; Sharma2022_pre sought to directly connect local structure to the local dynamics, in addition to investigating the global structure-dynamics correlation. For a fragile system, super-Arrhenius growth in the relaxation time implies a large change in the activation energy. It was suggested that the slope in the activation energy obtained from the macroscopic dynamics, when plotted as a function of the inverse depth of the caging potential, is high for more fragile liquids Nandi2021_prl . In a later study involving the machine learned softness parameter, it was shown that the slope of the activation energy obtained from microscopic dynamics, when plotted against the ML softness, increases with fragility Tah2022_jcp . A higher slope indicates that even small variations in the order parameter can lead to significant changes in the energy barrier and dynamics, reflecting a stronger coupling between structure and dynamics. Thus, these studies suggested that for a more fragile system, the structure-dynamics correlation is stronger, both macroscopically and microscopically.
@@
Our goal in this study is to understand how generic this statement is. To do that, we take three different sets of systems and vary fragility by changing different properties. We take a system interacting via the attractive Lennard Jones (LJ) potential and its repulsive counterpart, the Weeks-Chandler-Andersen(WCA) potential, and vary their fragility by changing the density Banerjee2016_jcp ; Berthier2009_jcp . We also vary the stepness of an attractive LJ-like potential and vary the fragility, which we refer to as the modified LJ $(q,p)$ system consisting of $(12,11)$, $(12,6)$ and $(8,5)$ models (where the pairs of integers refer to the exponent of the repulsive and attractive interactions)Sengupta2011_jcp ; Bordat2004_prl . This way, we have both attractive and repulsive potentials and also systems where the fragility is varied without changing the density.
@@
We find that there is no consistent correlation between fragility and the slope of the activation energy barrier. However, the latter does correlate well with a complementary measure of structure-dynamics correlation, as obtained from isoconfigurational ensemble simulations. Notably, two systems, LJ at $\rho=1.1$ and $(8,5)$, exhibit the highest structure-dynamics correlation, despite having very different fragility values. What is striking is that both systems also display a broad distribution of particle mobilities, a distinct double-peak in the logarithmic displacement distribution, and large values of the non-Gaussian parameter but small values of four-point dynamic susceptibility. We also find that these two systems are in the enthalpy-dominated regime and very close to the spinodal line where it is known that dynamic heterogeneity may arise due to mechanical instability and growing fluctuations Testard2014_jcp .
@@
These findings lead us to emphasise a critical distinction: dynamic heterogeneity is a broader concept than what is captured by $\chi_{4}$, which quantifies spatial correlations in particle motion. Dynamic heterogeneity also includes the distribution and variability of individual particle mobilities, which can arise due to different reasons and can exist even in the absence of significant spatial correlations. Thus, whereas $\chi_{4}$ highlights cooperative dynamics, other measures, such as the van Hove correlation function, the non-Gaussian parameter and mobility distributions from isoconfigurational runs—reflect heterogeneity in a more general sense.
@@
The rest of the paper is organized as follows. Section II describes the simulation details, while Section III outlines the methodology. Section IV presents the observations and results. Finally, Section V provides the conclusions. In addition, the paper includes four appendices that provide supplementary information.
@@ II.1 Simulation Details for the LJ and WCA Systems
$U_{ij}(r)=\begin{cases}\begin{aligned} &U_{ij}^{(LJ)}(r;\sigma_{ij},\epsilon_{ij})\\ &\quad-U_{ij}^{(LJ)}(r^{(c)}_{ij};\sigma_{ij},\epsilon_{ij}),\end{aligned}&r\leq r^{(c)}_{ij}\\ 0,&r>r^{(c)}_{ij},\end{cases}$ (1)
@@ II.2 Simulation Details for the $(q,p)$ Models
$U_{ij}(r)=\begin{cases}\begin{aligned} \epsilon_{ij}\left[\left(\dfrac{r^{\min}_{ij}}{r}\right)^{q}-\left(\dfrac{r^{\min}_{ij}}{r}\right)^{p}\right]\\ +c_{1,ij}r^{2}+c_{2,ij},\end{aligned}&r<r_{c}^{ij}\\ 0,&r\geq r_{c}^{ij}\end{cases}$ (2)
@@ III.1 Structural order parameter
To quantify the static structure in glassy systems, we analyze the Structural Order Parameter (SOP) at both macroscopic and microscopic levels. The SOP characterizes the depth of the mean-field caging potential experienced by a particle due to its surrounding neighbors and is derived from the Ramakrishnan-Yussouff free energy functional Ramakrishnan1979_prb . We define the macroscopic SOP as $(\beta\Phi)^{-1}$ and the microscopic SOP as $(\beta\phi)^{-1}$ throughout this study. The effective caging potential at zero displacement is computed as
[…]
@@ III.2 Entropy and local energy
$\begin{split}S^{i}_{u2}=-k_{B}\frac{\rho}{2}\sum_{v=1}^{2}x_{v}\int_{0}^{\infty}\left\{g^{i}_{uv}(r)\ln g^{i}_{uv}(r)-g^{i}_{uv}(r)+1\right\}\mathrm{d}r,\end{split}$ (8)
@@
$e^{i}_{uloc}=\frac{\rho}{2}\sum_{v=1}^{2}x_{v}\int_{0}^{\infty}U_{uv}(r)\,g^{i}_{uv}(r)\mathrm{d}r,$ (9)
@@ III.3 Relaxation time and Four-Point Susceptibility
The self-part of the van Hove correlation functionvanHove1954_review ; HansenMcDonald , which gives the probability distribution of single-particle displacements is written as:
@@
For systems in the high temperature regime, when the particle motion is homogeneous and diffusive, the van Hove correlation function is Gaussian. However, at the low temperature regime, due to heterogeneity in particle motion, the van Hove correlation function becomes non-Gaussian and it develops a tail which is best visualised in the probability distribution of logarithmic displacementsBarrat2010_soft , $P(\log_{10}\delta r)$, which is defined as:
$P(\log_{10}\delta r)=ln(10)4\pi\delta r^{3}G_{s}(\delta r,t).$ (14)
This representation emphasizes both mobile and immobile particle populations and is particularly useful for identifying the broad, often bimodal, displacement distributions observed at the $\alpha$-relaxation time in dynamically heterogeneous systems.
@@ IV RESULTS
As discussed in the Introduction, in this study we work with three different sets of systems, i) the attractive LJ potential at different densities ii) the repulsive WCA potential, at different densities Banerjee2016_jcp ; Berthier2009_jcp , iii) modified LJ $(q,p)$ sytem consisting of $(12,11)$,$(12,6)$ and $(8,5)$, systems- LJ-like potential where the steepness of the potential is varied Sengupta2011_jcp ; Bordat2004_prl , and vary fragility by changing different properties like the density and the steepness of the potential. While describing the results in all figures, we use a consistent color and symbol scheme to represent increasing fragility for each set: black circles for the least fragile systems, followed by red squares, green up-triangles, and blue down-triangles for progressively more fragile systems.
@@
IV.1 The temperature and SOP dependence of the energy barrier and its correlation with fragility
@@
We can describe the macroscopic energy barrier, $\Delta E^{ma}=T*ln(\tau_{\alpha}/\tau_{0})$, where $\tau_{\alpha}$ is the $\alpha$-relaxation time of the systems as obtained from the overlap function (Section III.3). $\tau_{0}$ is a prefactor. In Fig.1(left panels), we plot the macroscopic energy barrier as a function of inverse macroscopic SOP, $\beta\Phi$, for the three different sets of systems. We can also obtain the energy barrier from the microscopic motion of particles. This is a standard procedure Schoenholz2015_nature ; Sharma2022_pre , and the details are given in Appendix II. We can write the probability of a particle with a certain value of the inverse SOP to be a fast moving particle as, $P_{R}(\beta\phi)=P_{0}(\beta\phi)exp[-\Delta E(\beta\phi)/T]$. Here $\Delta E(\beta\phi)$ is the energy barrier for particles having the value of the inverse microscopic order parameter $\beta\phi$ (given in Eq.3). The microscopic energy barriers thus obtained are plotted against the inverse microscopic per particle level SOP, $\beta\phi$(right panels of Fig.2). We find that the rate of growth of both the macroscopic and the microscopic energy barriers show a similar trend. For the LJ and the WCA systems, we find that as the fragility of the systems increases, the average slope of the energy barrier increases. In a previous study involving some of us,Nandi2021_prl it was shown that the energy barrier as a function of the inverse SOP can be fitted in a straight line $\Delta E^{ma}=A+B*\beta\Phi$ where $A$ and $B$ are system-specific fitting parameters. This then suggests that the temperature dependence of the activation energy depends on the temperature dependence of the SOP. Thus, for strong liquids where the energy barrier is independent of temperature, it was shown that the parameter $B\rightarrow 0$ Nandi2021_prl and the fact that the slope $B$ increases with fragility is an expected result. A similar observation was also made in the ML studies, where it was shown that the slope of the microscopic energy barrier against the ML softness parameter increases with fragility Tah2022_jcp . However, the behaviour appears to be completely reversed when we consider the set of modified LJ, $(q,p)$ systems where the fragility is varied by changing the steepness of the potential. In this case, the slope of both the macroscopic and microscopic energy barriers is lowest for the $(8,5)$ system—the most fragile—and highest for the $(12,11)$ system—the least fragile. This finding challenges the commonly held assumption Nandi2021_prl ; Tah2022_jcp that the most fragile system should exhibit the steepest slope in the energy barrier versus the SOP, raising the question of whether this relationship is truly generic.
@@
Figure 1: Left Panels: Macroscopic energy barrier, $\Delta E^{\text{ma}}$, plotted against the inverse macroscopic structural order parameter, $\beta\Phi$, for three sets of systems: (a)LJ and (c)WCA, each examined at four different densities, and (e) modified LJ $(q,p)$ system. Right Panels: The inverse activation energy $1/\Delta E^{ma}$ plotted against $1/\Delta((\beta\phi)^{-1})$ for the same systems. The slope of the linear fits yields the $B^{\prime}/A^{\prime}$ values used in the fragility analysis(see Fig.4). The colour coding in the right panels matches that of the left panels.
@@
Note that this analysis depends on the order parameter we choose, and the fragility in terms of the order parameter can be different from that in terms of temperature. First, we present the temperature dependence of the SOP. Similar to earlier studies Nandi2021_prl ; Sharma2022_pre , we find that for all the systems the SOP, at macroscopic level, $(\beta\Phi)^{-1}$ and microscopic level, $<(\beta\phi)^{-1}>$ are linearly proportional to the temperature (right panel of Fig.2). Although the value of the order parameter at the macroscopic and microscopic levels is different, which is expected (as elaborated in Appendix I), their temperature dependences are similar. Both at the macroscopic and the microscopic levels, the SOP is linear with $T$ and can be fitted to the equations $(\beta\Phi)^{-1}=(\beta\Phi_{0})^{-1}+\gamma T$, and $<(\beta\phi)^{-1}>=<(\beta\phi_{0})^{-1}>+\gamma^{{}^{\prime}}T$ where $\gamma$ and $\gamma^{{}^{\prime}}$ are the corresponding slopes (given in Table 1). $(\beta\Phi_{0})^{-1}$ and $<(\beta\phi_{0})^{-1}>$ are the values of the macroscopic and average microscopic SOP, at T=0, respectively. We find that the values of the slopes $\gamma$ and $\gamma^{{}^{\prime}}$ are not identical but similar, which can be expected as both come from the temperature evolution of the radial distribution function.
@@
[…]
@@
Figure 2: Left Panels:Microscopic energy barrier, $\Delta E$, versus the inverse microscopic structural order parameter, $\beta\phi$, for three sets of systems: (b)LJ and (d)WCA, each examined at four different densities, and (f) modified LJ $(q,p)$ system. Right Panels:Temperature dependence of both the macroscopic structural order parameter, $(\beta\Phi)^{-1}$, (closed symbols) and the average microscopic structural order parameter $\langle(\beta\phi)^{-1}\rangle$,(open symbols) for the same systems. Dotted lines represent linear fits used to extract slope parameters for further analysis. The colour coding in the left panels matches that of the right panels.
@@
We next investigate the temperature and SOP dependence of the barrier $\Delta E^{ma}=T*ln(\tau_{\alpha}/\tau_{0})$ as predicted by the VFT expression. From the VFT expression, Eq.11 we can write,
$\Delta E^{ma}/T=\frac{T_{VFT}}{K_{VFT}(T-T_{VFT})}$ (18)
This clearly shows that the slope of $\Delta E^{ma}$ is not directly proportional but inversely proportional to the fragility. For both the LJ and WCA systems, fragility increases with density, accompanied by a corresponding rise in $T_{VFT}$ as the systems at higher densities operate in higher temperature regimes. This trend explains why we observe an increase in the slope, $\frac{T_{VFT}^{2}}{K_{VFT}}$, with increasing fragility. For the modified LJ $(q,p)$ system the fragility is higher for the $(8,5)$ system which has the lowest value of $T_{VFT}$ and this is the reason even in the $\Delta E^{ma}$ vs $1/T$ plots the $(8,5)$ system has the lowest slope (left panels of Fig.13 in the Appendix III). Interestingly the slope of $1/\Delta E^{ma}$ vs $1/T$ is the fragility (right panels of Fig.13 in Appendix III),
$1/\Delta E^{ma}=\frac{K_{VFT}}{T_{VFT}}-K_{VFT}.\frac{1}{T}$ (20)
We next investigate the SOP dependence of $\Delta E^{ma}$. In Eq.19 and Eq.20 if we replace $T$ in terms of SOP (Eq.LABEL:gamma) we can write,

### v4

@@ Abstract
Understanding the connection between structure, dynamics, and fragility, the rate at which the relaxation time grows with decreasing temperature, is central to unravelling the glass transition. Fragility is often associated with dynamic heterogeneity, implying that if structure influences dynamics, more fragile systems should exhibit stronger structure–dynamics correlations. In this study, we test the generality of this assumption using: Lennard-Jones (LJ) and Weeks–Chandler–Andersen (WCA) systems, where fragility is tuned via density, and a modified LJ $(q,p)$ system, where fragility is varied by changing the potential softness. We define a structural order parameter based on a mean-field caging potential and analyse energy barriers at both macroscopic and microscopic levels. While the macroscopic free energy barrier slope correlates with fragility, the microscopic free energy barrier does not show a consistent trend. Instead, it exhibits a strong correlation with a structure–dynamics correlation measure obtained from isoconfigurational ensemble simulations. Interestingly, the two systems showing the highest structure–dynamics correlation, LJ at $\rho=1.1$ and the $(8,5)$ model, are respectively the least and most fragile within their classes. These systems exhibit broad mobility distributions, large non-Gaussian parameters, yet low four-point susceptibilities, suggesting a decoupling between spatial correlation length and mobility contrast. Both systems lie in the enthalpy dominated regime and are close to the spinodal, pointing to mechanical instability as a source of heterogeneity. Our results reveal that structure–dynamics correlation is more closely linked to the contrast in individual particle mobility than to the spatial extent of dynamic correlations that typically scale with fragility.
@@ I Introduction
The behaviour of supercooled liquids near the glass transition is characterised by a dramatic slowing down in dynamics without any obvious structural transition Ediger2000_Physchem ; Debenedetti2001_Nature ; Dyre2006_modern ; Lubchenko2007_Physchem ; Tong2018_prx . The rapidity with which the relaxation time increases with decreasing temperature is quantified in terms of the fragility parameter Angell1995_science . Fragile liquids show a super-Arrhenius increase in relaxation time, while strong liquids follow an Arrhenius behaviour Martinez2001_nature .
@@
One of the hallmarks of supercooled liquids is dynamic heterogeneity, the presence of regions with significantly different mobilities. Traditionally, this heterogeneity has been linked to fragility: more fragile systems tend to exhibit stronger dynamic heterogeneity Kob1997_prl ; Donati1998_pre ; Sillescu1999_noncryst ; Ediger2000_Physchem ; Martinez2001_nature ; Sastry2001_nature ; Biroli2006_prl . This link is often assessed through quantities like the four-point dynamic susceptibility, $\chi_{4}$, which captures spatial correlations in particle dynamics Lacevic2003_jcp ; Toninelli2005_pre .
@@
Although, as discussed in the literature, the glass transition and the rapid growth of the dynamics happens without substantial structural change Debenedetti2001_Nature ; Lubchenko2007_Physchem , studies have shown that structure does indeed contribute to the dynamics, and small changes in structure can lead to large changes in other quantities like entropy which eventually strongly effects the dynamics Banerjee2014_prl ; Nandi2015_jcp ; Banerjee2016_jcp . There have also been studies trying to understand the structural origin of dynamical heterogeneity Cooper2008_nature ; Schoenholz2015_nature ; Richard2020_prm ; Nandi2021_prl . Among these studies, the machine learning study expressing the structural order by a softness parameter Schoenholz2015_nature and the microscopic study involving some of us expressing the structural order in terms of the depth of a mean field caging potential Nandi2021_prl ; Sharma2022_pre sought to directly connect local structure to the local dynamics, in addition to investigating the global structure-dynamics correlation. For a fragile system, super-Arrhenius growth in the relaxation time implies a large change in the activation energy. It was suggested that the slope in the activation energy obtained from the macroscopic dynamics, when plotted as a function of the inverse depth of the caging potential, is high for more fragile liquids Nandi2021_prl . In a later study involving the machine learned softness parameter, it was shown that the slope of the activation energy obtained from microscopic dynamics, when plotted against the ML softness, increases with fragility Tah2022_jcp . A higher slope indicates that even small variations in the order parameter can lead to significant changes in the free energy barrier and dynamics, reflecting a stronger coupling between structure and dynamics. Thus, these studies suggested that for a more fragile system, the structure-dynamics correlation is stronger, both macroscopically and microscopically.
@@
Our goal in this study is to understand how generic this statement is. To do that, we take three different sets of systems and vary fragility by changing different properties. We take a system interacting via the attractive Lennard Jones (LJ) potential and its repulsive counterpart, the Weeks-Chandler-Andersen(WCA) potential, and vary their fragility by changing the density Berthier2009_jcp ; Banerjee2016_jcp . We also vary the stepness of an attractive LJ-like potential and vary the fragility, which we refer to as the modified LJ $(q,p)$ system consisting of $(12,11)$, $(12,6)$ and $(8,5)$ models (where the pairs of integers refer to the exponent of the repulsive and attractive interactions)Bordat2004_prl ; Sengupta2011_jcp . This way, we have both attractive and repulsive potentials and also systems where the fragility is varied without changing the density.
@@
We find that there is no consistent correlation between fragility and the slope of the scaled activation free energy barrier. However, the latter does correlate well with a complementary measure of structure-dynamics correlation, as obtained from isoconfigurational ensemble simulations. Notably, two systems, LJ at $\rho=1.1$ and $(8,5)$, exhibit the highest structure-dynamics correlation, despite having very different fragility values. What is striking is that both systems also display a broad distribution of particle mobilities, a distinct double-peak in the logarithmic displacement distribution, and large values of the non-Gaussian parameter but small values of four-point dynamic susceptibility. We also find that these two systems are in the enthalpy-dominated regime and very close to the spinodal line Sastry2000_prl_press ; Sengupta2011_jcp where it is known that dynamic heterogeneity may arise due to mechanical instability and growing fluctuations Sastry2000_prl_press .
@@
These findings lead us to emphasise a critical distinction: dynamic heterogeneity is a broader concept than what is captured by $\chi_{4}$, which quantifies spatial correlations in particle motion. Dynamic heterogeneity also includes the distribution and variability of individual particle mobilities, which can arise due to different reasons and can exist even in the absence of significant spatial correlations. Thus, whereas $\chi_{4}$ highlights cooperative dynamics, other measures, such as the van Hove correlation function, the non-Gaussian parameter and mobility distributions from isoconfigurational runs, reflect heterogeneity in a more general sense.
@@
The rest of the paper is organised as follows. Section II describes the simulation details, while Section III outlines the methodology. Section IV presents the observations and results. Finally, Section V provides the conclusions. In addition, the paper includes five appendices that provide supplementary information.
@@ II.1 Simulation Details for the LJ and WCA Systems
[…]
@@ II.2 Simulation Details for the $(q,p)$ Models
$U_{ij}(r)=\begin{cases}\begin{aligned} \frac{\epsilon_{ij}}{q-p}\left[p\left(\dfrac{r^{\min}_{ij}}{r}\right)^{q}-q\left(\dfrac{r^{\min}_{ij}}{r}\right)^{p}\right]\\ +c_{1,ij}r^{2}+c_{2,ij},\end{aligned}&r<r_{c}^{ij}\\ 0,&r\geq r_{c}^{ij}\end{cases}$ (2)
@@ III.1 Structural order parameter
To quantify the static structure in glassy systems, we analyse the Structural Order Parameter (SOP) at both macroscopic and microscopic levels. The SOP is the inverse of the depth of the mean-field caging potential experienced by a particle due to its surrounding neighbours and is derived from the Ramakrishnan-Yussouff free energy functional Ramakrishnan1979_prb . We define the macroscopic SOP as $(\beta\Phi)^{-1}$ and the microscopic SOP as $(\beta\phi)^{-1}$ throughout this study. According to our earlier studies the inverse depth of the mean field caging potential is proportional to the curvature of the potential and thus related to its softness. Nandi2021_prl ; Sharma2022_pre The effective caging potential at zero displacement is computed as
$\begin{split}\beta\Phi&=-\rho\int\textbf{dr}\sum_{uv}C_{uv}(r)\,x_{u}x_{v}\,g_{uv}(r),\end{split}$ (3)
@@ III.2 Entropy and local energy
$\begin{split}S^{i}_{u2}=-k_{B}\frac{\rho}{2}\sum_{v=1}^{2}x_{v}\int_{0}^{\infty}\left\{g^{i}_{uv}(r)\ln g^{i}_{uv}(r)-g^{i}_{uv}(r)+1\right\}\textbf{dr},\end{split}$ (8)
@@
$e^{i}_{uloc}=\frac{\rho}{2}\sum_{v=1}^{2}x_{v}\int_{0}^{\infty}U_{uv}(r)\,g^{i}_{uv}(r)\textbf{dr},$ (9)
@@ III.3 Relaxation time and Four-Point Susceptibility
The self-part of the van Hove correlation functionvanHove1954_review ; HansenMcDonald , which gives the probability distribution of single-particle displacements, is written as:
@@
For systems in the high temperature regime, when the particle motion is homogeneous and diffusive, the van Hove correlation function is Gaussian. However, at the low temperature regime, due to heterogeneity in particle motion, the van Hove correlation function becomes non-Gaussian and it develops a tail which is best visualised in the probability distribution of logarithmic displacementsFlenner2005_pre ; Barrat2010_soft , $P(\log_{10}\delta r)$, which is related to the displacement distribution $P(\delta r)=4\pi\delta r^{2}G_{s}(r,t)$ and can be written as:
$P(\log_{10}\delta r)=ln(10)\delta rP(\delta r)=ln(10)4\pi\delta r^{3}G_{s}(\delta r,t).$ (14)
This representation emphasises both mobile and immobile particle populations and is particularly useful for identifying the broad, often bimodal, displacement distributions observed at the $\alpha$-relaxation time in dynamically heterogeneous systems.
@@ IV RESULTS
As discussed in the Introduction, in this study we work with three different sets of systems, i) the attractive LJ potential at different densities ii) the repulsive WCA potential, at different densities Banerjee2016_jcp ; Berthier2009_jcp , iii) modified LJ $(q,p)$ sytem consisting of (12,11)(12,11),$(12,6)$ and $(8,5)$, systems- LJ-like potential where the steepness of the potential is varied Sengupta2011_jcp ; Bordat2004_prl , and vary fragility by changing different properties like the density and the steepness of the potential. While describing the results in all figures, we use a consistent colour and symbol scheme to represent increasing fragility for each set: black circles for the least fragile systems, followed by red squares, green up-triangles, and blue down-triangles for progressively more fragile systems.
@@
IV.1 The temperature and SOP dependence of the free energy barrier and its correlation with fragility
@@
We can describe the macroscopic free energy barrier, $\Delta E^{ma}=T*ln(\tau_{\alpha}/\tau_{0})$, where $\tau_{\alpha}$ is the $\alpha$-relaxation time of the systems as obtained from the overlap function (Section III.3). $\tau_{0}$ is a prefactor. In Fig.1(left panels), we plot the macroscopic free energy barrier as a function of inverse macroscopic SOP, $\beta\Phi$, for the three different sets of systems. We can also obtain the free energy barrier from the microscopic motion of particles. This is a standard procedure Schoenholz2015_nature ; Sharma2022_pre , and the details are given in Appendix II. We can write the probability of a particle with a certain value of the inverse SOP to be a fast moving particle as, $P_{R}(\beta\phi)=P_{0}(\beta\phi)exp[-\Delta E(\beta\phi)/T]$. Here $\Delta E(\beta\phi)$ is the free energy barrier for particles having the value of the inverse microscopic order parameter $\beta\phi$ (given in Eq.3). The microscopic free energy barriers thus obtained are plotted against the inverse microscopic per particle level SOP, $\beta\phi$(left panels of Fig.2). We observe that the macroscopic free energy barrier grows more strongly than linear, approximately quadratic, while the microscopic free energy barrier increases linearly. This weaker growth of the microscopic barrier arises from its inability to capture cooperative molecular motion, consistent with earlier findings Tah2022_jcp . As we will show later, the stronger than linear growth of the macroscopic free energy barrier as a function SOP can be derived from the VFT expression, which clearly shows that the prediction of the linear growth is an approximation. Nevertheless, both macroscopic and microscopic free energy barriers exhibit similar qualitative trends. In both LJ and WCA systems, we find that the free energy barrier grows more steeply with increasing fragility, in agreement with previous studies Nandi2021_prl ; Tah2022_jcp . However, the behaviour appears to be completely reversed when we consider the set of modified LJ, $(q,p)$ systems where the fragility is varied by changing the steepness of the potential. In this case, the slope of both the macroscopic and microscopic free energy barriers is lowest for the $(8,5)$ system, the most fragile, and highest for the $(12,11)$ system, the least fragile. This finding challenges the commonly held assumption Nandi2021_prl ; Tah2022_jcp that the most fragile system should exhibit the steepest slope in the free energy barrier versus the SOP, raising the question of whether this relationship is truly generic.
@@
Figure 1: Left Panels: Macroscopic free energy barrier, $\Delta E^{\text{ma}}$, plotted against the inverse macroscopic structural order parameter, $\beta\Phi$, for three sets of systems: (a)LJ and (c)WCA, each examined at four different densities, and (e) modified LJ $(q,p)$ system. Right Panels: The inverse activation energy $1/\Delta E^{ma}$ plotted against $1/\Delta((\beta\phi)^{-1})$ for the same systems. The slope of the linear fits yields the $B^{\prime}/A^{\prime}$ values used in the fragility analysis(see Fig.4). The colour coding in the right panels matches that of the left panels.
@@
Note that this analysis depends on the order parameter we choose, and the fragility in terms of the order parameter can be different from that in terms of temperature. First, we present the temperature dependence of the SOP. Similar to earlier studies Nandi2021_prl ; Sharma2022_pre , we find that for all the systems the SOP, at macroscopic level, $(\beta\Phi)^{-1}$ and microscopic level, $(\beta\phi)^{-1}>$ are linearly proportional to the temperature (right panels of Fig.2). Although the value of the order parameter at the macroscopic and microscopic levels is different, which is expected (as elaborated in Appendix I), their temperature dependences are similar. Both at the macroscopic and the microscopic levels, the SOP is linear with $T$ and can be fitted to the equations $(\beta\Phi)^{-1}=(\beta\Phi_{0})^{-1}+\gamma T$, and $(\beta\phi)^{-1}>=<(\beta\phi_{0})^{-1}>+\gamma^{{}^{\prime}}T$ where $\gamma$ and $\gamma^{{}^{\prime}}$ are the corresponding slopes (given in Table 1). $(\beta\Phi_{0})^{-1}$ and $(\beta\phi_{0})^{-1}>$ are the values of the macroscopic and average microscopic SOP, at T=0, respectively. We find that the values of the slopes $\gamma$ and $\gamma^{{}^{\prime}}$ are not identical but similar, which can be expected as both come from the temperature evolution of the radial distribution function.
@@
We next investigate the fragility of the systems both as a function of temperature, given by Eq.11, and also as a function of $\Delta((\beta\Phi)^{-1})$, which can be represented by the following VFT expression GarciaColin1989_vft ,
@@
Figure 2: Left Panels:Microscopic free energy barrier, $\Delta E$, versus the inverse microscopic structural order parameter, $\beta\phi$, for three sets of systems: (b)LJ and (d)WCA, each examined at four different densities, and (f) modified LJ $(q,p)$ system. Right Panels:Temperature dependence of both the macroscopic structural order parameter, $(\beta\Phi)^{-1}$, (closed symbols) and the average microscopic structural order parameter $\langle(\beta\phi)^{-1}\rangle$,(open symbols) for the same systems. Dotted lines represent linear fits used to extract slope parameters for further analysis. The colour coding in the left panels matches that of the right panels.
@@
If we now assume that $\Delta E^{ma}=T*ln(\tau_{\alpha}/\tau_{0})$ then using Eq.17 we can write an expression for $\Delta E^{ma}$ in terms of the SOP where we replace the temperature from Eq.LABEL:gamma. We can then write,
$1/\Delta E^{ma}=\frac{K_{VFT}}{T_{VFT}}-\gamma.K_{VFT}.\frac{1}{\Delta((\beta\Phi)^{-1})}=A^{\prime}-B^{\prime}.\frac{1}{\Delta((\beta\Phi)^{-1})}$ (18)
and